OpenStudy (anonymous):

use the rational root theorem to list all possible rational roots of the polynomial equation x3-x2-8x+=4 a. -4,-1,1,4 b.-4,-2,-1,1,2,4 c.no roots d.1,2,4

OpenStudy (anonymous):

+=4?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

or what do you mean

OpenStudy (anonymous):

What does plus equal mean here? Is our equation\[x^3 -x^2-8x = 4 \]or\[x^3 -x^2-8x+(something) = 4 \]

OpenStudy (anonymous):

it is x3-x2-8x+4=0

OpenStudy (anonymous):

Ok, then our rational roots will be the all factors of 4 over the all factors of 1. The 1 comes from the coefficient of x^3.

OpenStudy (anonymous):

'possible

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so would it be the awnser b

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thank you hey can you help me on another problem

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

find the roots of x3-2x2+5x-10=0

OpenStudy (anonymous):

a.2+isquare root 5 b. 1 +5 c.5 + square root 7 d. -1 + square root 3

OpenStudy (anonymous):

Apply the Rational Root Test. So, the possible roots are plus or minus 1, 2, 5, 10. We can test these to see if they are solutions to the equation.

OpenStudy (anonymous):

Does option b need to be rewritten?